Science and medicine. A passenger train can travel 325 mi in the same time a freight train takes to travel 200 mi. If the speed of the passenger train is 25 mi/h faster than the speed of the freight train, find the speed of each.
Answer by stanbon(48558)

Science and medicine. A passenger train can travel 325 mi in the same time a freight train takes to travel 200 mi. If the speed of the passenger train is 25 mi/h faster than the speed of the freight train, find the speed of each.
-----------
Passenger train DATA:
distance=325 mi ; Rate = x+25 mph; Time = d/r = 325/(x+25)
----------
Freight train DATA:
distance = 200 mi ; Rate = x mph; Time = d/r = 200/x
--------------
EQUATION:
time = time
325/(x+25) = 200/x
325x = 200x+500
125x = 500
x=4 mph for freight train
x+25 = 29 mph for passenger train
==================
Cheers,
Stan H.